OpenStudy (anonymous):

What is the standard deviation of the following data? If necessary, round your answer to two decimal places. 6, 7, 10, 11, 11, 13, 16, 18, 25 5.48 5.58 5.68 5.78

OpenStudy (unklerhaukus):

\[\begin{matrix}{} j &|& 6 &7 & 10 & 11&13&16&18&25 \\ \hline N(j)&|& 1 &1& 1 & 2 & 1 & 1&1&1 \end{matrix}\]

OpenStudy (unklerhaukus):

\[N=\sum\limits_{j=0} ^{\infty} N(j)\]

OpenStudy (anonymous):

5.58

OpenStudy (unklerhaukus):

\[\langle j\rangle=\sum\limits_{j=0}^{\infty} jP(j)=\sum\limits_{j=0}^{\infty} j\frac{N(j)}{N}\]\[=\frac{1}{N}\sum\limits_{j=0}^{\infty} j{N(j)}\]

OpenStudy (anonymous):

is my answer correct?

OpenStudy (unklerhaukus):

i dont know i haven't got that far yet, what did you do?

OpenStudy (unklerhaukus):

\[\langle j^2 \rangle =\sum\limits_{j=0}^{\infty}j^2P(j)=\frac{1}{N}\sum\limits_{j=0}^{\infty}j^2N(j)\]

OpenStudy (unklerhaukus):

\[\sigma^2={\langle(\Delta j)^2\rangle} \] \[\sigma=\sqrt{\langle(\Delta j)^2\rangle} =\sqrt{\langle j^2 \rangle -\langle j \rangle^2}\]

OpenStudy (unklerhaukus):

for N i am getting 9,

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Data: 6,7,10,11[2],13,16,18,25, n= 9 Σx= 117 , Σx²= 1801 , µ= 13 , σn-1=5.91608, σn=5.577734

OpenStudy (anonymous):

?

OpenStudy (unklerhaukus):

for the average im am getting 13

OpenStudy (unklerhaukus):

average squared i am getting 169 the square of the average im am getting 1447/9~164.111

OpenStudy (anonymous):

ok

OpenStudy (unklerhaukus):

so the difference (ie the standard deviation) is 4.888,.... damn i messed up

OpenStudy (unklerhaukus):

your Σx= 117, µ= 13 are correct, check your Σx²= 1801 i am getting 1477

OpenStudy (unklerhaukus):

I think i am doing this wrong somehow however you had σn=5.577734 which matches with b. so that is probably it